# Gauges and Accelerated Optimization over Smooth and/or Strongly Convex Sets

### v3

@@ 1 Introduction
Critically, our proposed algorithms are projection-free. The typical first-order method assumption of being able to project onto $S_{i}$ limits algorithms to simple constraints. Instead, we only assume oracles for one-dimensional linesearches (to evaluate the gauges defined in (5)) and to compute normal vectors on the boundary of each $S_{i}$. In contrast, projected gradient methods require an orthogonal projection oracle. (Projections onto sets parallel computing often costly proximal operators of functions, whereas computing normal vectors parallels cheaper gradient calculations.) For example, if $S_{i}$ is an ellipsoid, orthogonal projection lacks a closed form, but our linesearch and its normal vectors have closed forms, costing a single matrix-vector multiplication to compute. For polyhedrons, projection is a quadratic program, whereas a normal vector is computed by identifying any one active constraint.
@@ 1.1 Related Works
where $\hat{\iota}_{S}(x)$ is the nonstandard indicator function $\hat{\iota}_{S}(x)=\begin{cases}+\infty&\text{ if }x\in S\\ 0&\text{ if }x{\color[rgb]{0.0,0.0,0.0}\notin}S\end{cases}$ can be equivalently solved by minimizing the radially dual problem
@@
where $f^{\Gamma}(y)=\sup\{v>0\mid vf(y/v)\leq 1\}$ [1, Proposition 24, Proposition 11]. Note one can have $f(0)>0$ and hence a positive maximum value without loss of generality by translating the problem to place a known feasible point $x_{0}\in\mathrm{int\ }\left(S\cap\{x\mid f(x)\in\mathbb{R}\}\right)$ to the origin and adding a constant. This reformulation is unconstrained and only depends on the constraints $S$ via their gauge. Consequently, our structural results on gauges will facilitate the direct use of accelerated optimization methods on this radial dual (squared). Prior radial methods have only used simple subgradient methods [12, 13] or smoothing techniques [14, 2].
@@ 2 Preliminaries on Smoothness and Strong Convexity
For a closed convex differentiable function $h$, the following are equivalent:
@@
    $h(y)\leq h(x)+\nabla h(x)^{T}(y-x))+\frac{\tilde{L}}{2}\|y-x\|^{2}\ .$
@@
For a closed convex function $h$, the following are equivalent:
@@
We abuse notation, saying any potentially nonsmooth set is $\infty$-smooth, viewing the limit above as reducing to $\bar{y}\in S$. In the other limit, the only 00-smooth sets are half-spaces. Mirroring Proposition 1, several equivalent characterizations of smooth sets are given below, proof deferred to Appendix A.2.3111 Characterizations of more general nonconvex notions of smooth sets are given in [25, Definition 3.2, Page 286]. However, we are unable to find any statement and proof of such equivalent characterizations for simply convex sets based on convex normal vectors. Hence our proofs of these characterizations may be of some independent interest..
@@
For a closed convex set $S$, the following are equivalent:
@@
Note (b) implies Lipschitzness of unit normal vectors $\|\zeta_{1}-\zeta_{2}\|\leq\beta\|x_{1}-x_{2}\|$. The above condition (d), which we refer to as being (locally) $\beta$-smooth w.r.t. $(\bar{y},\zeta)$, does not imply the other conditions: For any $\epsilon>0$, $S=\{(x,y)\mid|x|\leq\epsilon\}$ is globally only $1/\epsilon$-smooth but locally 00-smooth. We expect this implication to hold among compact sets but are unaware of a proof of this result.
@@
In particular, this reduces to saying $S$ is 00-strongly convex if $S$ is convex. In the other limit, singletons are the only $\infty$-strongly convex sets. Mirroring Proposition 2, below are equivalent characterizations, proven in Appendix A.2.4.
@@
For a closed convex bounded set $S$, the following are equivalent:
@@ 2.1 Common Families of Smooth/Strongly Convex Sets
Figure 1: Examples of strongly convex and smooth set with the inner approximation (3) in red and the outer approximation (4) in blue. The $p=1.5$-norm ball in (a) is strongly convex but not smooth. The $p=2$-norm ellipsoid in (b) is smooth and strongly convex. The $p=3$-norm ball in (c) is smooth but not strongly convex.
@@
Note $\frac{g}{\|g\|}$ is a unit length normal vector of $S$ at $x$. Then $L$-smoothness ensures $h(y)\leq z$ holds whenever $h(x)+g^{T}(y-x)+\frac{L}{2}\|y-x\|^{2}\leq z$ (i.e., $y\in B(x-\frac{\|\nabla h(x)\|}{L}\frac{g}{\|g\|},\frac{\|g\|}{L})$). Similarly, $\mu$-strong convexity ensures $h(y)\leq z$ implies $h(x)+g^{T}(y-x)+\frac{\mu}{2}\|y-x\|^{2}\leq z$ (i.e., $y\in B(x-\frac{\|g\|}{\mu}\frac{g}{\|g\|},\frac{\|g\|}{\mu})$). ∎
@@
Consider any closed convex $h$ that is locally $L$-smooth, $\mu$-strongly convex w.r.t. some $x$ and $g\in\partial h(x)$ and let $(\zeta,\delta)=(g,-1)/\|(g,-1)\|$. Then epi $\text{epi }h=\{(x,t)\mid h(x)\leq t\}$ is $L\delta$-smooth and $\mu\delta^{3}$-strongly convex w.r.t. $((x,h(x)),(\zeta,\delta))$.
This implies the epigraph of any $L$-smooth function is globally $L$-smooth since $\delta\leq 1$. No similar global result holds for strongly convex functions since globally strongly convex sets must be bounded.
@@
Consider any closed, $\beta$-smooth, $\alpha$-strongly convex set $S$ and linear $A\colon\mathcal{E}\rightarrow\mathcal{E}^{\prime}$, $b\in\mathcal{E}^{\prime}$ with full column rank. Then the set $\{x\mid Ax-b\in S\}$ is $\beta\frac{\lambda_{\max}(A^{T}A)}{\sqrt{\lambda_{\min}(A^{T}A)}}$-smooth and $\alpha\frac{\lambda_{\min}(A^{T}A)}{\sqrt{\lambda_{\max}(A^{T}A)}}$-strongly convex.
@@
Without loss of generality, let $b=0$ and consider $A^{+}=(A^{T}A)^{-1}A^{T}$. As an intermediate result, we first claim $A^{+}B(0,1/\beta)$ is $\beta\frac{\lambda_{\max}(A^{T}A)}{\sqrt{\lambda_{\min}(A^{T}A)}}$-smooth. Observe that
$A^{+}B(0,1/\beta)=\{y\mid\|Ay\|^{2}\leq 1/\beta^{2}\}\ .$
The inclusion $\subseteq$ holds as for any $A^{+}x\in A^{+}B(0,1/\beta)$, $y=A^{+}x$ has $\|Ay\|=\|AA^{+}x\|\leq\|AA^{+}\|\|x\|\leq 1/\beta$ since $AA^{+}$ is an idempotent matrix with eigenvalues 0 and 1. The inclusion $\supseteq$ holds as for any $y$ with $\|Ay\|\leq 1/\beta$, $x=Ay\in B(0,1/\beta)$ and so $y=A^{+}x\in A^{+}B(0,1/\beta)$. Then Lemma 6 with $h(y)=\|Ay\|^{2}$ proves this claim.
Since $S$ is $\beta$-smooth, by Proposition 3(c), there exists a closed convex set $S_{0}$, such that $S_{0}+B(0,\frac{1}{\beta})=S$. In particular, $A^{+}S_{0}+A^{+}B(0,\frac{1}{\beta})=A^{+}S$. Proposition 3(c) applied to $A^{+}B(0,\frac{1}{\beta})$ further gives a closed convex set $B_{0}$ such that $B_{0}+B\left(0,\frac{1}{\beta}\frac{\sqrt{\lambda_{\min}(A^{T}A)}}{\lambda_{\max}(A^{T}A)}\right)=A^{+}B(0,\frac{1}{\beta})$. Hence $\left(A^{+}S_{0}+B_{0}\right)+B\left(0,\frac{1}{\beta}\frac{\sqrt{\lambda_{\min}(A^{T}A)}}{\lambda_{\max}(A^{T}A)}\right)=A^{+}S$ and so applying Proposition 3(c) again with the closed convex set $A^{+}S_{0}+B_{0}$, we conclude $A^{+}S$ is $\beta\frac{\lambda_{\max}(A^{T}A)}{\sqrt{\lambda_{\min}(A^{T}A)}}$-smooth. Symmetric reasoning shows $S$ is $\alpha\frac{\lambda_{\min}(A^{T}A)}{\sqrt{\lambda_{\max}(A^{T}A)}}$-strongly convex. ∎
@@ 3 The Structure of Gauges of Structured Sets
Consider any $y\in\mathcal{E}$ and closed convex set $S$ with $0\in\mathrm{int\ }S$. If $\gamma_{S}(y)\neq 0$ and at the point $\bar{y}=y/\gamma_{S}(y)$, $S$ is $\alpha$-strongly convex w.r.t. $\left(\bar{y},\zeta\right)$ for some unit normal vector $\zeta\in N_{S}(\bar{y})$, then $\frac{1}{2}\gamma_{S}^{2}$ is strongly convex with parameter
$\frac{1}{2(\zeta^{T}\bar{y})^{3}}\left(\zeta^{T}\bar{y}+\alpha\|\bar{y}\|^{2}-\sqrt{(\zeta^{T}\bar{y}+\alpha\|\bar{y}\|^{2})^{2}-4\alpha(\zeta^{T}\bar{y})^{3}}\right)$
w.r.t. $(y,g)$, where $g=\frac{\zeta}{\zeta^{T}\bar{y}}\in\partial(\frac{1}{2}\gamma_{S}^{2})(y)$. If $\gamma_{S}(y)=0$ and $y=0$, then $\{0\}=\partial(\frac{1}{2}\gamma_{S}^{2})(y)$ and $\frac{1}{2}\gamma_{S}^{2}$ is $\frac{1}{D(S)^{2}}$-strong convex w.r.t. $(0,0)$. If $\gamma_{S}(y)=0$ and $y\neq 0$, $\frac{1}{2}\gamma_{S}^{2}(y)$ is not strongly convex at $y$.
Note that this quantity can be lower bounded by the following $O(\alpha)$ quantity:
@@
Consider any $y\in\mathcal{E}$ and closed convex set $S$ with $0\in\mathrm{int\ }S$. If $\gamma_{S}(y)\neq 0$ and at the point $\bar{y}=y/\gamma_{S}(y)$, $S$ is $\beta$-smooth w.r.t. $\left(\bar{y},\zeta\right)$ for some unit normal vector $\zeta\in N_{S}(\bar{y})$, then $\frac{1}{2}\gamma_{S}^{2}$ is smooth with parameter
$\frac{1}{2(\zeta^{T}\bar{y})^{3}}\left(\zeta^{T}\bar{y}+\beta\|\bar{y}\|^{2}+\sqrt{(\zeta^{T}\bar{y}+\beta\|\bar{y}\|^{2})^{2}-4\beta(\zeta^{T}\bar{y})^{3}}\right)$
w.r.t. $(y,g)$, where $g=\nabla(\frac{1}{2}\gamma_{S}^{2})(y)=\frac{\zeta}{\zeta^{T}\bar{y}}$. If $\gamma_{S}(y)=0$, then $\{0\}=\partial(\frac{1}{2}\gamma_{S}^{2})(y)$ and $\frac{1}{2}\gamma_{S}^{2}$ is $\frac{1}{R(S)^{2}}$-smooth w.r.t. $(y,0)$.
Note that this quantity can be upper bounded by the following $O(\beta)$ quantity:
@@
The above two theorems provide the gauge squared’s local strongly convex/smooth parameter. Then the gauge square’s global strongly convex/smooth parameter can be obtained by bounding these over all boundary points $\bar{y}$ due to Propositions 1 and 2. Noting $\|\bar{y}\|\leq D(S)$ and $\zeta^{T}\bar{y}\geq R(S)$ by convexity gives the following global bounds.
@@ 3.1 Examples
These theorems/corollaries immediately establish the structure for the gauge squared of many common families of constraints. Here we discuss a few such examples, namely, halfspaces =$\mathcal{H}=\{x\mid a^{T}x\leq b\}$, $p$-norm unit balls $B_{p}=\{x\mid\|x\|_{p}\leq 1\}$, and $p$-norm ellipsoids $E_{p}=\{x\mid\|Ax-b\|_{p}\leq 1\}$. Table 1 summarizes these results, showing the smoothness and strong convexity of each $S$ and its half gauge squared. These examples will also be utilized in our numerical evaluations in Section 5.
@@
$B_{p}$ $p\in(1,2)$ $S$ $(p-1)n^{\frac{1}{2}-\frac{1}{p}}$infty
$\frac{1}{2}\gamma^{2}_{S}$ $(p-1)n^{\frac{1}{2}-\frac{1}{p}}$infty
@@
$E_{p}$ $p\in(1,2)$ $S$ $\frac{\lambda_{\min}}{\sqrt{\lambda_{\max}}}(p-1)n^{\frac{1}{2}-\frac{1}{p}}$infty
$\frac{1}{2}\gamma^{2}_{S}$ $C^{SC}_{p,b}\lambda_{\min}(p-1)n^{\frac{1}{2}-\frac{1}{p}}$infty
@@
$\frac{1}{2}\gamma^{2}_{S}$ $\frac{\lambda_{\min}}{(1+\|b\|_{2})(2+\|b\|_{2})}$ $\frac{\lambda_{\max}(2-\|b\|_{2})}{(1-\|b\|_{2})^{2}}$
@@
Table 1: Strongly convexity and smoothness constants. For the $p$-norm ellipsoids, $\lambda_{\max}$ and $\lambda_{\min}$ denote the maximum and minimum eigenvalues of $A^{T}A$, $C^{SC}_{p,b}=\frac{1}{(1+\|b\|_{2})(2+\|b\|_{2})}$, and $C^{SM}_{p,b}=\frac{9}{2}\left(\frac{p}{1-\|b\|_{p}^{p}}\right)^{3}$.
Halfspaces. Consider any halfspace with =$\mathcal{H}=\{x\mid a^{T}x\leq b\}$. Such sets are not strongly convex ($\alpha=0$) but are infinitely smooth ($\beta=0$) everywhere. To ensure the gauge is well-defined, we require $0\in\mathrm{int\ }\mathcal{H}$ (that is, $b>0$). Note $D(\mathcal{H})=\infty$ and $R(\mathcal{H})=b/\|a\|_{2}$. Consider any $\bar{y}$ on the boundary of $\mathcal{H}$ and unit normal $\zeta=a/\|a\|_{2}\in N_{\mathcal{H}}(\bar{y})$. Note $\zeta^{T}\bar{y}=b/\|a\|_{2}$. Then Theorem 11 vacuously implies $\mu=0$-strong convexity and Theorem 12 implies $L=\|a\|_{2}^{2}/b^{2}$-smoothness at $\bar{y}$.
@@
$p$-Norm Balls. Consider any $p\in(1,\infty)$-norm unit ball $B_{p}=\{x\mid\|x\|_{p}\leq 1\}$. Depending on $p$, this ball and its gauge squared are either smooth or strongly convex. Namely [3, Lemma 4] have shown $B_{p}$ and $\frac{1}{2}\gamma_{B_{p}}^{2}$ are both $\alpha=\mu=(p-1)n^{\frac{1}{2}-\frac{1}{p}}$-strongly convex whenever $p\in(1,2]$ and are $\beta=L=(p-1)n^{\frac{1}{2}-\frac{1}{p}}$-smooth whenever $p\in[2,\infty)$. Note that neither of these bounds are tight for $p\neq 2$. Our theory could be applied to yield tighter, although substantially less elegant, bounds. The details of such an approach are given in Appendix A.1.
$p$-Norm Ellipsoids. As our last example, we consider generalizing the example above to ellipsoidal sets $E_{p}=\{x\mid\|Ax-b\|_{p}\leq 1\}$ for any $p\in(1,\infty)$. Note for the gauge to be well-defined (i.e., having $0\in E_{p}$), we require $\|b\|_{p}\leq 1$. The following lemma, mirroring Lemma 8, allows us to bound the strong convexity and smoothness of such a set’s gauge squared.
@@
If a set $S$ has $\frac{1}{2}\gamma_{S}^{2}$ $\alpha$-strongly convex or $\beta$-smooth, then $E=\{x\mid x\in AS\}$ has $\frac{1}{2}\gamma^{2}_{E}(y)=\frac{1}{2}\gamma^{2}_{S}(Ay)$ being $\lambda_{\min}(A^{T}A)\alpha$-strongly convex or $\lambda_{\max}(A^{T}A)\beta$-smooth, respectively.

### v4

@@ 1 Introduction
Critically, our proposed algorithms are projection-free. The typical first-order method assumption of being able to project onto $S_{i}$ limits algorithms to simple constraints. Instead, we only assume oracles for one-dimensional linesearches (to evaluate the gauges defined in (5)) and to compute normal vectors on the boundary of each $S_{i}$. In contrast, projected gradient methods require an orthogonal projection oracle. (Computing projections onto sets is comparable to computing proximal operators of functions, which is often costly, whereas computing normal vectors is comparable to cheaper gradient calculations.) For example, if $S_{i}$ is an ellipsoid, orthogonal projection lacks a closed form, but our linesearch and its normal vectors have closed forms, costing a single matrix-vector multiplication to compute. For polyhedrons, projection is a quadratic program, whereas a normal vector is computed by identifying any one active constraint.
@@ 1.1 Related Works
where $\hat{\iota}_{S}(x)$ is the nonstandard indicator function $\hat{\iota}_{S}(x)=\begin{cases}+\infty&\text{ if }x\in S\\ 0&\text{ if }x\notin S\end{cases}$ can be equivalently solved by minimizing the radially dual problem
@@
where $f^{\Gamma}(y)=\sup\{v>0\mid vf(y/v)\leq 1\}$ [1, Proposition 24]. Note one can have $f(0)>0$ and hence a positive maximum value without loss of generality by translating the problem to place a known feasible point $x_{0}\in\mathrm{int\ }\left(S\cap\{x\mid f(x)\in\mathbb{R}\}\right)$ to the origin and adding a constant. This reformulation is unconstrained and only depends on the constraints $S$ via their gauge. Consequently, our structural results on gauges will facilitate the direct use of accelerated optimization methods on this radial dual (squared). Prior radial methods have only used simple subgradient methods [12, 13] or smoothing techniques [14, 2].
@@ 2 Preliminaries on Smoothness and Strong Convexity
For a closed convex differentiable function $h$ and a constant $L\in(0,\infty)$, the following are equivalent:
@@
    $h(y)\leq h(x)+\nabla h(x)^{T}(y-x)+\frac{\tilde{L}}{2}\|y-x\|^{2}\ .$
@@
For a closed convex function $h$ and a constant $\mu\in(0,\infty)$, the following are equivalent:
@@
We abuse notation, saying any potentially nonsmooth set is $\infty$-smooth, viewing the limit above as reducing to $\bar{y}\in S$. In the other limit, the only 0-smooth sets are half-spaces. Mirroring Proposition 1, several equivalent characterizations of smooth sets are given below, proof deferred to Appendix A.2.3111Characterizations of more general nonconvex notions of smooth sets are given in [25, Definition 3.2, Page 286]. However, we are unable to find any statement and proof of such equivalent characterizations for simply convex sets based on convex normal vectors. Hence our proofs of these characterizations may be of some independent interest..
@@
For a closed convex set $S$ and a constant $\beta\in(0,\infty)$, the following are equivalent:
@@
Note (b) implies Lipschitzness of unit normal vectors $\|\zeta_{1}-\zeta_{2}\|\leq\beta\|x_{1}-x_{2}\|$. The above condition (d), which we refer to as being (locally) $\beta$-smooth w.r.t. $(\bar{y},\zeta)$, does not imply the other conditions: For any $\epsilon>0$, $S=\{(x,y)\mid|x|\leq\epsilon\}$ is globally only $1/\epsilon$-smooth but locally 0-smooth. We expect this implication to hold among compact sets but are unaware of a proof of this result.
@@
In particular, this reduces to saying $S$ is 0-strongly convex if $S$ is convex. In the other limit, singletons are the only $\infty$-strongly convex sets. Mirroring Proposition 2, below are equivalent characterizations, proven by the equivalences of $\mathbf{(e),(h),(f),(l)}$ in [25, Theorem 2.1, page 265-266].
@@
For a closed convex bounded set $S$ and a constant $\alpha\in(0,\infty)$, the following are equivalent:
@@ 2.1 Common Families of Smooth/Strongly Convex Sets
Figure 1: Examples of strongly convex and smooth sets with the inner approximation (3) in red and the outer approximation (4) in blue. The $p=1.5$-norm ball in (a) is strongly convex but not smooth. The $p=2$-norm ellipsoid in (b) is smooth and strongly convex. The $p=3$-norm ball in (c) is smooth but not strongly convex.
@@
Note $\frac{g}{\|g\|}$ is a unit length normal vector of $S$ at $x$. Then $L$-smoothness ensures $h(y)\leq z$ holds whenever $h(x)+g^{T}(y-x)+\frac{L}{2}\|y-x\|^{2}\leq z$ (i.e., $y\in B(x-\frac{\|g\|}{L}\frac{g}{\|g\|},\frac{\|g\|}{L})$). Similarly, $\mu$-strong convexity ensures $h(y)\leq z$ implies $h(x)+g^{T}(y-x)+\frac{\mu}{2}\|y-x\|^{2}\leq z$ (i.e., $y\in B(x-\frac{\|g\|}{\mu}\frac{g}{\|g\|},\frac{\|g\|}{\mu})$). ∎
@@
Consider any closed convex $h$ that is locally $L$-smooth, $\mu$-strongly convex w.r.t. some $x$ and $g\in\partial h(x)$ and let $(\zeta,\delta)=(g,-1)/\|(g,-1)\|$. Then $\text{epi }h=\{(x,t)\mid h(x)\leq t\}$ is locally $L|\delta|$-smooth and $\mu|\delta|^{3}$-strongly convex w.r.t. $((x,h(x)),(\zeta,\delta))$.
This implies the epigraph of any $L$-smooth function is globally $L$-smooth since $|\delta|\leq 1$. No similar global result holds for strongly convex functions since globally strongly convex sets must be bounded.
@@
Consider any closed, $\beta$-smooth, $\alpha$-strongly convex set $S$ and linear invertible $A\colon\mathcal{E}\rightarrow\mathcal{E}^{\prime}$, $b\in\mathcal{E}^{\prime}$. Then the set $\{x\mid Ax-b\in S\}$ is $\beta\frac{\lambda_{\max}(A^{T}A)}{\sqrt{\lambda_{\min}(A^{T}A)}}$-smooth and $\alpha\frac{\lambda_{\min}(A^{T}A)}{\sqrt{\lambda_{\max}(A^{T}A)}}$-strongly convex.
@@
Without loss of generality, consider $b=0$. Note $\{x\mid Ax\in S\}=A^{-1}S$. Since $S$ is $\beta$-smooth, by Proposition 3(c), there exists a closed convex set $S_{0}$, such that $S_{0}+B(0,\frac{1}{\beta})=S$. In particular, $A^{-1}S_{0}+A^{-1}B(0,\frac{1}{\beta})=A^{-1}S$. Proposition 3(c) applied to $A^{-1}B(0,\frac{1}{\beta})$ further gives a closed convex set $B_{0}$ such that $B_{0}+B\left(0,\frac{1}{\beta}\frac{\sqrt{\lambda_{\min}(A^{T}A)}}{\lambda_{\max}(A^{T}A)}\right)=A^{-1}B(0,\frac{1}{\beta})$. Since $A^{-1}B(0,\frac{1}{\beta})$ is a compact ellipsoid, $B_{0}$ has to be a compact set. It follows $A^{-1}S_{0}+B_{0}$ is closed and convex. Hence $\left(A^{-1}S_{0}+B_{0}\right)+B\left(0,\frac{1}{\beta}\frac{\sqrt{\lambda_{\min}(A^{T}A)}}{\lambda_{\max}(A^{T}A)}\right)=A^{-1}S$ and so applying Proposition 3(c) again with the closed convex set $A^{-1}S_{0}+B_{0}$, we conclude $A^{-1}S$ is $\beta\frac{\lambda_{\max}(A^{T}A)}{\sqrt{\lambda_{\min}(A^{T}A)}}$-smooth. Symmetric reasoning shows $S$ is $\alpha\frac{\lambda_{\min}(A^{T}A)}{\sqrt{\lambda_{\max}(A^{T}A)}}$-strongly convex. ∎
@@ 3 The Structure of Gauges of Structured Sets
Consider some constant $\alpha\in(0,\infty)$, any $y\in\mathcal{E}$ and closed convex set $S$ with $0\in\mathrm{int\ }S$. If $\gamma_{S}(y)\neq 0$ and at the point $\bar{y}=y/\gamma_{S}(y)$, $S$ is $\alpha$-strongly convex w.r.t. $\left(\bar{y},\zeta\right)$ for some unit normal vector $\zeta\in N_{S}(\bar{y})$, then $\frac{1}{2}\gamma_{S}^{2}$ is strongly convex with parameter
$\frac{1}{2(\zeta^{T}\bar{y})^{3}}\left(\zeta^{T}\bar{y}+\alpha\|\bar{y}\|^{2}-\sqrt{(\zeta^{T}\bar{y}+\alpha\|\bar{y}\|^{2})^{2}-4\alpha(\zeta^{T}\bar{y})^{3}}\right)$ (17)
w.r.t. $(y,g)$, where $g=\frac{\gamma_{S}(y)\zeta}{\zeta^{T}\bar{y}}\in\partial(\frac{1}{2}\gamma_{S}^{2})(y)$. If $\gamma_{S}(y)=0$ and $y=0$, then $\{0\}=\partial(\frac{1}{2}\gamma_{S}^{2})(y)$ and $\frac{1}{2}\gamma_{S}^{2}$ is $\frac{1}{D(S)^{2}}$-strong convex w.r.t. $(0,0)$. If $\gamma_{S}(y)=0$ and $y\neq 0$, $\frac{1}{2}\gamma_{S}^{2}(y)$ is not strongly convex at $y$.
Note that (17) can be lower bounded by the following $O(\alpha)$ quantity:
@@
Consider some constant $\beta\in(0,\infty)$, any $y\in\mathcal{E}$ and closed convex set $S$ with $0\in\mathrm{int\ }S$. If $\gamma_{S}(y)\neq 0$ and at the point $\bar{y}=y/\gamma_{S}(y)$, $S$ is $\beta$-smooth w.r.t. $\left(\bar{y},\zeta\right)$ for some unit normal vector $\zeta\in N_{S}(\bar{y})$, then $\frac{1}{2}\gamma_{S}^{2}$ is smooth with parameter
$\frac{1}{2(\zeta^{T}\bar{y})^{3}}\left(\zeta^{T}\bar{y}+\beta\|\bar{y}\|^{2}+\sqrt{(\zeta^{T}\bar{y}+\beta\|\bar{y}\|^{2})^{2}-4\beta(\zeta^{T}\bar{y})^{3}}\right)$ (19)
w.r.t. $(y,g)$, where $g=\nabla(\frac{1}{2}\gamma_{S}^{2})(y)=\frac{\gamma_{S}(y)\zeta}{\zeta^{T}\bar{y}}$. If $\gamma_{S}(y)=0$, then $\{0\}=\partial(\frac{1}{2}\gamma_{S}^{2})(y)$ and $\frac{1}{2}\gamma_{S}^{2}$ is $\frac{1}{R(S)^{2}}$-smooth w.r.t. $(y,0)$.
Note that (19) can be upper bounded by the following $O(\beta)$ quantity:
@@
The above two theorems provide the gauge squared’s local strongly convex/smooth parameter. Then the gauge square’s global strongly convex/smooth parameter can be obtained by bounding these over all possible points $\bar{y}$ (i.e. all boundary points) due to Propositions 1 and 2. We claim $\|\bar{y}\|\leq D(S)$ and $\zeta^{T}\bar{y}\geq R(S)$. The first claim is immediate from the definition of $D(S)$. The second follows by considering the support function as $\zeta^{T}\bar{y}=\sup_{z\in S}\{\zeta^{T}z\}\geq\sup_{z\in B(0,R(S))}\zeta^{T}z=R(S)$. We then have the following global bounds.
@@ 3.1 Examples
These theorems/corollaries above immediately establish the structure for the gauge squared of many common families of constraints. Here we discuss a few such examples, namely, halfspaces $\mathcal{H}=\{x\mid a^{T}x\leq b\}$, $p$-norm unit balls $B_{p}=\{x\mid\|x\|_{p}\leq 1\}$, and $p$-norm ellipsoids $E_{p}=\{x\mid\|Ax-b\|_{p}\leq 1\}$. Table 1 summarizes these results, showing the smoothness and strong convexity of each $S$ and its half gauge squared. These examples will also be utilized in our numerical evaluations in Section 5.
@@
$B_{p}$ $p\in(1,2)$ $S$ $(p-1)n^{\frac{1}{2}-\frac{1}{p}}$ $\infty$
$\frac{1}{2}\gamma^{2}_{S}$ $(p-1)n^{\frac{1}{2}-\frac{1}{p}}$ $\infty$
@@
$E_{p}$ $p\in(1,2)$ $S$ $\frac{\lambda_{\min}}{\sqrt{\lambda_{\max}}}(p-1)n^{\frac{1}{2}-\frac{1}{p}}$ $\infty$
$\frac{1}{2}\gamma^{2}_{S}$ $C^{SC}_{p,b}\lambda_{\min}(p-1)n^{\frac{1}{2}-\frac{1}{p}}$ $\infty$
@@
$\frac{1}{2}\gamma^{2}_{S}$ $\frac{\lambda_{\min}}{(1+\|b\|_{2})+(1+\|b\|_{2})^{2}}$ $\frac{\lambda_{\max}(2-\|b\|_{2})}{(1-\|b\|_{2})^{2}}$
@@
Table 1: Strong convexity and smoothness constants. For the $p$-norm ellipsoids, $\lambda_{\max}$ and $\lambda_{\min}$ denote the maximum and minimum eigenvalues of $A^{T}A$, $C^{SC}_{p,b}=\frac{1}{(1+\|b\|_{2})+(p-1)n^{\frac{1}{2}-\frac{1}{p}}(1+\|b\|_{2})^{2}}$, and $C^{SM}_{p,b}=\frac{9}{2}\left(\frac{p}{1-\|b\|_{p}^{p}}\right)^{3}$.
Halfspaces. Consider any halfspace with $\mathcal{H}=\{x\mid a^{T}x\leq b\}$. Such sets are not strongly convex ($\alpha=0$) but are infinitely smooth ($\beta=0$) everywhere. To ensure the gauge is well-defined, we require $0\in\mathrm{int\ }\mathcal{H}$ (that is, $b>0$). Note $D(\mathcal{H})=\infty$ and $R(\mathcal{H})=b/\|a\|_{2}$. Consider any $\bar{y}$ on the boundary of $\mathcal{H}$ and unit normal $\zeta=a/\|a\|_{2}\in N_{\mathcal{H}}(\bar{y})$. Note $\zeta^{T}\bar{y}=b/\|a\|_{2}$. Then Theorem 11 vacuously implies $\mu=0$-strong convexity and Theorem 12 implies $L=\|a\|_{2}^{2}/b^{2}$-smoothness at $\bar{y}$.
@@
$p$-Norm Balls. Consider any $p\in(1,\infty)$-norm unit ball $B_{p}=\{x\mid\|x\|_{p}\leq 1\}$. Depending on $p$, this ball and its gauge squared are either smooth or strongly convex. Namely [3, Lemma 4] showed $B_{p}$ and $\frac{1}{2}\gamma_{B_{p}}^{2}$ are both $\alpha=\mu=(p-1)n^{\frac{1}{2}-\frac{1}{p}}$-strongly convex whenever $p\in(1,2]$ and are $\beta=L=(p-1)n^{\frac{1}{2}-\frac{1}{p}}$-smooth whenever $p\in[2,\infty)$. Note that neither of these bounds are tight for $p\neq 2$. Our theory could be applied to yield tighter, although substantially less elegant, bounds. The details of such an approach are given in Appendix A.1.
$p$-Norm Ellipsoids. As our last example, we consider generalizing the example above to ellipsoidal sets $E_{p}=\{x\mid\|Ax-b\|_{p}\leq 1\}$ for any $p\in(1,\infty)$ and $A$ invertible. Note for the gauge to be well-defined (i.e., having $0\in E_{p}$), we require $\|b\|_{p}\leq 1$. The following lemma, mirroring Lemma 8, allows us to bound the strong convexity and smoothness of such a set’s gauge squared.
@@
If a set $S$ has $\frac{1}{2}\gamma_{S}^{2}$ $\alpha$-strongly convex or $\beta$-smooth, then $E=\{x\mid Ax\in S\}$ with invertible $A$, has $\frac{1}{2}\gamma^{2}_{E}(y)=\frac{1}{2}\gamma^{2}_{S}(Ay)$ being $\lambda_{\min}(A^{T}A)\alpha$-strongly convex or $\lambda_{\max}(A^{T}A)\beta$-smooth, respectively.